40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 207 of 341  Not logged in ELOG logo
ID Date Author Type Category Subject
  6750   Mon Jun 4 23:48:43 2012 jenneUpdateGreen Lockinglowered gain

We're trying to do a yarm measurement....before I forget, I want to write this down...

I changed the gain of each of the top 2 SR560's down, by a factor of 2.  This made the overload lights quit coming on.

  6749   Mon Jun 4 17:14:31 2012 JenneUpdateLSCLSC recompiled several times today

As of now, the regular LSC DoF triggers work, just as they used to.  There is a problem with the filter module triggers that I haven't figured out yet. 

We can't send integers (like control words for the filter banks) through Choice blocks, since those pass doubles by default.  I fixed that by removing the choice block, but the triggering still isn't happening properly.

  6748   Sun Jun 3 23:50:00 2012 DenUpdateCDSbiquad=1

From now all models calculate iir filters using biquad form. I've added biquad=1 to cdsParameters to all models except c1cal, built, installed and restarted them.

  6747   Sun Jun 3 01:30:07 2012 DenUpdatePEMsts-2 and guralp in isolation box

We have 2 sts-2 readout box - pink and blue. Pink outputs 12 DVC - this a problem of amplifier. This box has a rectifier (the box works from AC power) and an amplifier for velocity channels. Mass positions, calibration channels are connected by a wire from input to output. The amplifier for velocity channels does not work properly, so I connected velocity channels directly to the output - the signal from sts-2 is large enough even without amplification. When I plugged sts-2 to pink readout board, on the velocity output I saw ~4 VDC. Sts-2 was needed to be recentered. I pressed AUTOZERO command, but this did not work out. Before I had checked that this readout box indeed gives an autozero logical signal - 5VDC for ~2 sec. I think it does not provides sts-2  with enough current, seismometer needs 0.1 A in autozero regime.

Blue readout box after switching it to 1 sec regime and zeroing sts-2 started to output reasonable signal for gains = 10. I tried gains = 100, X velocity channel started to output noise. Now the gain is 10 and the response is 120 sec. But at least this box works. Still performance is not clear as well as noise level. To determine this I've put sts-2 to isolation box.

DSC_4315.JPG                    DSC_4319.JPG

After I've put Guralps in the isolation and waited for a couple of days, Guralp noise has been improved a little more.

lin_box_noise.png                 mcl_gur.png

  6746   Sat Jun 2 03:19:37 2012 yutaUpdateGreen LockingY green beat note found? - too small

Summary:
  I tried to find Y arm green beat in order to do the mode scan.
  I found a beat peak(see attached picture), but the amplitude seems too small.
  It is may be because the alignment/mode matching of the green beams at the PSL table is so bad. Or, the peak I found might be a beat from junk light.

What I did:
  1. Aligned Y arm to the IR beam from MC.

  2. Re-aligned Y end green beam to the Y arm using steering mirrors on the Y end table.

  3. Re-aligned PSL green optics.

  # C1:GCV-GREEN_TRY is temporary connected to the DC output of the Y green beat PD.

  4. Temperature of the PSL laser was 31.48 deg C, so I set "T+" of the Y end laser to 34.47 deg C, according to Bryan's formula (elog #4439);

  Y_arm_Temp_set = 0.87326*T_PSL + 6.9825

  5. Scanned Y end laser temperature by C1:GCY-SLOW_SERVO2_OFFSET. Starting value was 29725 and I scanned from 27515 to 31805, by 10 or 100. Laser frequency changes ~ 6 MHz / 10 counts, so it means that I scanned ~ 2.5 GHz. During the scan, I toggled C1:AUX-GREEN_Y_Shutter to make sure the green beam resonates in TEM00 mode.

  # I made a revolutionary python script for toggling channels(/opt/rtcds/caltech/c1/scripts/general/toggler.py). I made it executable.

  6. Found a tiny beat note when C1:GCY-SLOW_SERVO2_OFFSET = 29815. I confirmed it is a beat signal by blocking each PSL and Y arm green beam into the beat PD. I left  C1:GCY-SLOW_SERVO2_OFFSET = 29815.

  7. I found that Bryan's formula;

Y_arm_Temp_meas = 0.95152*T_PSL + 3.8672
Y_arm_Temp_set = 0.87326*T_PSL + 6.9825

  was actually

Y_arm_Temp_set = 0.95152*T_PSL + 3.8672
Y_arm_Temp_meas = 0.87326*T_PSL + 6.9825

  according to his graph(elog #4439). So, I set  "T+" of the Y end laser to 33.82 deg C.

  8. This time, I scanned PSL laser temperature by C1:PSL-FSS_SLOWDC. I found a tiny beat note when C1:PSL-FSS_SLOWDC = 1.0995. C1:PSL-FSS_SLOWDC has 10 V range, so I scanned ~ 10 GHz, assuming the laser frequency changes 1 GHz/K and the temperature changes 1 K/V.

  9. Re-aligned PSL green optics so that the beam hits optics at their center, and checked that the poralization of the two green beams are the same.

  10. Checked that amplifier ZFL-100LN+ on the beat PD is working correctly. The power was supplied correctly (+15 V) and measured gain was ~ 25 dBm.

  11. Exchanged BNC cable which connects the beat PD to the spectrum analyzer. Previous one we used was too long and it had -15 dB loss(measured). I exchanged to shorter one which has -2 dB loss.

Beat note amplitude estimation:
  The amplitude of the beat note observed in the spectrum analyzer was ~ -54 dBm. According to the estimation below, it seems too small.

  The measured power of the two green beams are

  P_Y = 4 uW
  P_PSL = 90 uW

  So, the power of the beat signal should be

  P_beat ~ 2 sqrt(P_Y * P_PSL) = 37 uW

  Responsivity and transimpedance of the beat PD (Broadband PD, LIGO-T0900582) are 0.3 A/W and 2 kOhm. So, the power of the electrical signal is

  W = (P_beat * 0.3 A/W * 2 kOhm / sqrt(2))^2 / 50 Ohm = 5 uW

  5 uW is -23 dBm. We have +25 dB amplifier after the PD and the loss of the BNC cable is -2 dB. So, if the two beams interfere perfectly, the peak height of the beat signal should be ~ 0 dBm. The measured value -54 dBm seems too small. According to elog #5860, measured value by Kiwamu and Katrin was -36 dBm.

Current values:
  PSL laser temperature: 31.48 deg C (PSL HEPA 100%)
  Y end laser "T+": 33.821 deg C
  Y end laser "ADJ": 0
  C1:GCY-SLOW_SERVO2_OFFSET = 29815 (was 29725)

Attachment 1: CIMG1437.JPG
CIMG1437.JPG
  6745   Fri Jun 1 19:48:54 2012 ranaUpdateSUSOplevs all different, messed up

 

 Its a good question, but the answer is yes. At some level all of the OL servos should be different to handle the different mechanical resonances of the suspension as well as the optical table's acoustic noise and the different noise requirements for the difference optical cavities.

However, it would be better to have the same basic structure and then one or two customization filters.

  6744   Fri Jun 1 18:07:32 2012 steveUpdateSUSOplevs servo values

Quote:

For some reason the state of the oplevs is completely different for almost every suspension.  They have different sets of filters in the bank, and different filters engaged.  wtf?  How did this happen?  Is this correct?  Do we expect that the state of the oplevs should be different on all the different suspensions?  I wouldn't have thought so.

I discovered this because the PRM is unstable with the oplevs engaged.  I don't think it was yesterday.  Is something hidden changing the oplev settings?

 As of February 23, 2012 when oplev PIT and YAW transfer functions were taken

 

OPLEV SERVO 300^ 2:0 BR ELP RLP RES GAIN QPD counts OD mm           
                   
ETMY pit 300^ 2:0 BR 35 80 0.5 -1.5 15,500 ~1  
ETMY yaw 300^ 2:0 BR 35 80 0.6 -1.0      
ETMX pit 300^ 2:0 BR 35 80 0.5 0.5 1,500 ~1.5  
ETMX yaw 300^ 2:0 BR 35 80 0.6 1.0      
ITMY pit 300^ 2:0 BR   80   2.0 15,000 ~2.5  
ITMY yaw 300^ 2:0 BR   80   -4.0      
ITMX pit 300^ 2:0 BR   80   1.0 1,350 ~1.5  
ITMX yaw 300^ 2:0 BR   80   -2.0      
BS pit 300^ 2:0 BR 50     0.5 3,500 ~1  
BS yaw 300^ 2:0   50   3.3 -1.0      
PRM pit 300^ 2:0 BR     3.3 1.0 4,000 ~2  
PRM yaw 300^ 2:0 BR 35   3.3,  4 -0.5      
SRM pit 300^ 2:0 BR 40     -2.0 2,600 ~2.5  
SRM yaw 300^ 2:0 BR 40     2.0      

 

  6743   Fri Jun 1 14:56:08 2012 JamieUpdateSUSOplevs all different, messed up

For some reason the state of the oplevs is completely different for almost every suspension.  They have different sets of filters in the bank, and different filters engaged.  wtf?  How did this happen?  Is this correct?  Do we expect that the state of the oplevs should be different on all the different suspensions?  I wouldn't have thought so.

I discovered this because the PRM is unstable with the oplevs engaged.  I don't think it was yesterday.  Is something hidden changing the oplev settings?

Attachment 1: oplevs.png
oplevs.png
  6742   Fri Jun 1 14:40:24 2012 JamieUpdateComputersc1sus and c1iscex - bad fb connections

Quote:

This is at least partially related to the mx_stream issue I reported previously.  I restarted mx_stream on c1iscex and that cleared up the models on that machine.

Something else is happening with c1sus.  Restarting mx_stream on c1sus didn't help.  I'll try to fix it when I get over there later.

I managed to recover c1sus.  It required stopping all the models, and the restarting them one-by-one:

$ rtcds stop all     # <-- this does the right to stop all the models with the IOP stopped last, so they will all unload properly.

$ rtcds start iop

$ rtcds start c1sus c1mcs c1rfm

I have no idea why the c1sus models got wedged, or why restarting them in this way fixed the issue.

  6741   Fri Jun 1 10:00:03 2012 steveOmnistructureGeneral40m Wireless Network

Quote:

Mike Pedraza came by today to install a new wireless network router configured for the 40m lab network.  It has a 'secret' SSID i.e. not meant for public use outside the lab.  You can look up the password and network name on the rack.  Pictures below show the location of the labels.

40m_Wifi.png

 Mike P swapped in a new network router Linksys E1000

  6740   Fri Jun 1 09:50:50 2012 JamieUpdateComputersc1sus and c1iscex - bad fb connections

Quote:

Something bad happened to c1sus and c1iscex ~20 min ago.  They both have "0x2bad" 's.  I restarted the daqd on the framebuilder, and then rebooted c1sus, and nothing changed.  The SUS screens are all zeros (the gains seem to be set correctly, but all of the signals are 0's).

If it's not fixed when I get in tomorrow, I'll keep poking at it to make it better.

 This is at least partially related to the mx_stream issue I reported previously.  I restarted mx_stream on c1iscex and that cleared up the models on that machine.

Something else is happening with c1sus.  Restarting mx_stream on c1sus didn't help.  I'll try to fix it when I get over there later.

  6739   Fri Jun 1 08:17:47 2012 steveUpdateIOOPMC trends

Quote:

IOO Angle & IOO Position QPDs centered.

 PMC trend of 400 and 1200 days

The Innolight 2W based PSL- IOO was implemented in the ~ summer of 2010

Attachment 1: PMC400d.png
PMC400d.png
Attachment 2: PMC1200d.png
PMC1200d.png
  6738   Fri Jun 1 08:01:46 2012 steveUpdateComputersc1sus and c1iscex are down

Quote:

Something bad happened to c1sus and c1iscex ~20 min ago.  They both have "0x2bad" 's.  I restarted the daqd on the framebuilder, and then rebooted c1sus, and nothing changed.  The SUS screens are all zeros (the gains seem to be set correctly, but all of the signals are 0's).

If it's not fixed when I get in tomorrow, I'll keep poking at it to make it better.

 

 

Attachment 1: compdown.png
compdown.png
  6737   Fri Jun 1 02:33:40 2012 JenneUpdateComputersc1sus and c1iscex - bad fb connections

Something bad happened to c1sus and c1iscex ~20 min ago.  They both have "0x2bad" 's.  I restarted the daqd on the framebuilder, and then rebooted c1sus, and nothing changed.  The SUS screens are all zeros (the gains seem to be set correctly, but all of the signals are 0's).

If it's not fixed when I get in tomorrow, I'll keep poking at it to make it better.

  6736   Fri Jun 1 02:13:00 2012 JenneUpdateGreen LockingAttempt at Ygreen beat - failed

[Yuta, Jenne]

We tried to find the Ygreen beat note, with no success yet.  We calculate from Bryan's formula that the Yend laser should be ~34.68C.  But Katrin has an elog saying that she was looking around 19C.  I don't know why the discrepancy, but maybe this is part of our problem?  Kiwamu elog-responded that the epics output had to be high (~9V) when the temp was 19C.  So maybe we need a smaller offset setting in the slow servo with the 34C temperature?

We set the "T+" on the Ygreen laser controller to 34.68C using the dial, and then tried a few large steps with the offset in the Ygreen slow servo.  The idea was to see if we could swing past the beat, so we would know vaguely where it was.  But we never saw a resonance on the spectrum analyzer, even with a "hold max" trace. 

We confirmed that there is signal going to the SLOW input of the laser controller's front panel.  Yuta watched a voltmeter while I changed the epics value, and we successfully changed the signal.  However, after plugging the SLOW cable back in, we noticed that no matter what we set the epics value to, we don't see any temperature change reported on the front panel display.  There is something in the manual( according to Katrin) that the "LT" display is not accurate when a cable is plugged in.  But none of the display values changed.  I think there is a measured temp output on the back that Bryan mentioned that we could use to see if something is really changing inside.

Anyhow, no beatnote found yet tonight.  We confirmed before starting that the alignment onto the beat PD was good, so that's not the problem.

  6735   Thu May 31 23:53:00 2012 JenneUpdateLSCLSC trigger update

I modified the lsc model (after Jamie finished) to use a new triggering scheme.  It HAS NOT yet been compiled and tested, since it's way past time for us to start beatnote-ing.  I will compile, test, debug, etc. tomorrow. Don't compile the LSC model tonight. 

Now we also have (assuming no bugs.....) triggering capability for the filter modules in the filter banks.  Yay!  Testing, etc will commence tomorrow.

  6734   Thu May 31 22:13:08 2012 JamieUpdateCDSc1lsc: added remaining SHMEM senders for ERR and CTRL, c1oaf model updated appropriately

All the ERR and CTRL outputs in c1lsc now go to SHMEM senders.  I renamed the the CTRL output SHMEM senders to be more generic, since they aren't specifically for OAF anymore.  See attached image from c1lsc.

c1oaf was updated so that SHMEM receivers pointed to the newly renamed senders.

c1lsc and c1oaf were rebuilt, installed, and restarted and are now running.

Attachment 1: lsc-shmem-out.png
lsc-shmem-out.png
  6733   Thu May 31 17:47:25 2012 SureshOmnistructureGeneral40m Wireless Network

Mike Pedraza came by today to install a new wireless network router configured for the 40m lab network.  It has a 'secret' SSID i.e. not meant for public use outside the lab.  You can look up the password and network name on the rack.  Pictures below show the location of the labels.

40m_Wifi.png

  6732   Thu May 31 16:54:12 2012 JenneUpdateGreen LockingLinks to old elogs for green beatnote laser temps

Because I keep taking a long time to search for these, since I can't remember the keywords in the different entries, here are the links:

elog 3759 : Green X end aux laser temperature setting vs. PSL laser temperature setting

elog 4439 : Green Y end aux laser temperature setting vs. PSL laser temperature setting

More words: beat note, doubling, second harmonic.

Relevant results: 

T_Xend = 8.31 + 0.9293*T_PSL

T_Yend = 6.9825 + 0.87326*T_PSL

 

Also, C1:GCY-SLOW_SERVO2_OFFSET was 29725 (twenty nine thousand seven hundred twenty five) cts when we sat down to start today.

C1:GCX-SLOW_SERVO2_OFFSET was 80 (eighty) cts when we sat down to start today.  Why the offsets are so different, I don't know.  But I was able to find the X green beatnote with this small number offset, so it is approximately correct.

  6731   Thu May 31 16:19:07 2012 yutaUpdateGreen Lockingtemperature setting for PSL doubling crystal

I fixed the temperature control of the oven for the PSL doubling crystal.
The PID settings were not good, and also, TC200 was beging DETUNED. So, I activated TUNE function and adjusted PID settings.
I'm not sure what the DETUNE function is for. The manual can be found here;
   http://www.thorlabs.com/thorproduct.cfm?partnumber=TC200

Current settings for Thorlabs TC200 are (Red ones are what I changed from the previous setting);

parameters Xend Yend PSL
TEMP SET (deg C) 37.5 35.7 36.9
P 250 250 250
I 60 60 200 (was 117)
D 25 25 40 (was 19)
(DE)TUNE on? TUNE TUNE TUNE (was DETUNE)
TMAX (deg C) 200 200 170
PMAX (Watts) 18 18 18
temperature sensor PTC100 PTC100 PTC100
  6730   Thu May 31 11:38:19 2012 DenUpdatePEMisolation system

I've put Guralps into the Steve's 2 box isolation system. Noise got better, coherence between 2 seismometers improved. We still need better performance. Probably, one device is noisy and we can not determine which one using these 2 seismometers. We need more seismometers. Sadly, STS-2 readout box is not working.

lin_box_noise2.png    lin_box_coh.png

  6729   Thu May 31 11:02:14 2012 steveUpdateLockingETMX 1064 trans camera

Quote:

Quote:

...will be helpful for acquiring lock after the vent.  We should install a camera at ETMX.

 Do that.

 Jenne and Steve

Sony CCD in place needs alignment and ND filter

  6728   Thu May 31 10:31:19 2012 JamieUpdateIOOMC beam spot oscillation

Quote:

This is a common occurrence when diagnostic scripts are written without the ability to handle exceptions (e.g. ctrl-c, terminal gets closed, etc.).

The first thing to do is make sure that the "new" script you are writing doesn't already exist (hint: look in the old scripts directory).

If you are writing a script that touches things in the interferometer, it must always return the settings to the initial state on abnormal termination:

http://linuxdevcenter.com/pub/a/linux/lpt/44_12.html

This is very good advice.  However, "trap" is bash-specific.  tcsh has a different method that uses a function called "onint".  Here's a description of the difference.

A couple notes about bash traps:

  • You can give a name instead of a number for the signal.  So instead of trap 'do stuff' 1 you can say trap 'do stuff' SIGHUP
  • The easiest signal to use is EXIT, which covers all your bases (ie. anything that would cause the script to exit prematurely.
  • You can define a function that gets executed in the trap

So the easiest way to use it is something like the following:

#!/bin/bash   # define cleanup function  function cleanup {      # do cleanup stuff, like reset EPICS records to defaults      ....  }  # set the trap on EXIT  trap cleanup EXIT  # the rest of your script below here
...
  6727   Thu May 31 04:03:17 2012 yutaUpdateIOOscript for MC beam spot measurement

I wrote a wrapping script for measuring MC beam spot. We had to run several scripts for the measurement (see elog #6688), but now, you only need to run /opt/rtcds/caltech/c1/scripts/ASS/MC/mcassMCdecenter.

The measured data file will be stored in /opt/rtcds/caltech/c1/scripts/ASS/MC/dataMCdecenter/ directory, with a timestamp.
The calculated beam spot position data will be logged in /opt/rtcds/caltech/c1/scripts/ASS/MC/dataMCdecenter/logMCdecenter.txt file.
I had to edit sensemcass.m file, in order to write the result into the log file. In this way, we can keep track of the beam displacement.

Currently, the calculation script is written in the MATLAB file(sensemcass.m), which isn't very nice.
To run a MATLAB file from the command line
, you have to write something like this;

matlab -nodesktop -nosplash -r "sensemcass('./dataMCdecenter/MCdecenter201205210258.dat')"

 

  6726   Thu May 31 02:27:24 2012 yutaUpdateIOOscript for reliefing MC WFS

I wrote a simple script for reliefing MC WFS servo. The script is located at /opt/rtcds/caltech/c1/scripts/MC/reliefMCWFS.
It simply uses ezcaservo to minimize the offset of the WFS feedback signal using MC alignment sliders.

    ezcaservo -r C1:SUS-MC${optic}_ASC${dof}_OUT16 -s 0 -g 0.0001 -t 10 C1:SUS-MC${optic}_${dof}_COMM


I put "MC WFS relief" button on the WFS medm screen (/opt/rtcds/caltech/c1/medm/c1ioo/master/C1IOO_WFS_MASTER.adl).

  6725   Thu May 31 01:36:17 2012 yutaUpdateGreen LockingGREEN_TRX/GREEN_TRY PDs

I did the cabling for monitoring DC transmission of the green beam from the end table.
The two PDs are called GREEN TRX and GREEN TRY, and the channel names are C1:GCV-GREEN_TRX and C1:GCV-GREEN_TRY.
The two signal from the PDs go to the ADC_0 card of the c1ioo computer.

Now, C1:GCV-GREEN_TRX/Y are actually connected to the respective PDs, but green beams are not hitting on the PD. We need two pickoff mirrors.

  6724   Thu May 31 01:27:16 2012 yutaUpdateGreen LockingPSL and Y arm green beams aligned

[Jenne, Yuta]

We aligned the PSL green optics so that the PSL green beam and Y arm green beam interfere. 2 beams are now hitting the Y arm beat PD. The DC level from the beat PD is about 13 mV.

We didn't try to see the beat signal for today, because the temperature of the doubling crystal seemed funny. We need to look into it tommorow.

Currently, the temperature control is enabled and the set point is 36.9 deg C, but the temperature is stuck at 33.0 deg C.

  6723   Thu May 31 01:20:41 2012 JenneUpdateASSASS filter outputs are non-zero with no input

I was looking a little at ASS, while Yuta was doing some Green transmitted DC PD work, and I find that the output of some filters is totally insane with no deliberate input or excitation signals.

Note in the figure that the filter (which is a 2nd order butter bandpass in the C1:ASS-LOCKIN29_SIG filter bank) is ringing a lot - this needs fixing.  But, more disconcertingly, sometimes (not every time) the arm flashes, the input to the filter bank gets a ~1 sample long spike that is ~9,000,000 counts.  9 million is a lot of counts.  This is then making the filter go crazy.

Any ideas on how this can happen, and how we can stop / fix it?  It's certainly a CDS issue, but I'm not sure where or how.

Attachment 1: ASS_Yarm_crazyOutputs_30May2012.png
ASS_Yarm_crazyOutputs_30May2012.png
  6722   Thu May 31 00:56:13 2012 JamieMetaphysicsComputersPlease remember to check in code changes

I know it's really hard to remember, but our future selves will thank us dearly if we remember to commit all of our code changes to the svn with nice log messages.  At the moment there's a LOT of modified stuff in the userapps working directory that needs to be committed:

controls@pianosa:/opt/rtcds/userapps/release 0$ svn status | grep '^M'
M       cds/c1/models/c1rfm.mdl
M       sus/c1/medm/templates/SUS_SINGLE.adl
M       sus/c1/models/c1mcs.mdl
M       sus/c1/models/c1sus.mdl
M       sus/c1/models/c1scx.mdl
M       sus/c1/models/c1scy.mdl
M       isc/c1/models/c1pem.mdl
M       isc/c1/models/c1ioo.mdl
M       isc/c1/models/ADAPT_XFCODE_MCL.c
M       isc/c1/models/c1oaf.mdl
M       isc/c1/models/c1gcv.mdl
M       isc/common/medm/OAF_OVERVIEW.adl
M       isc/common/medm/OAF_DOF_BLRMS.adl
M       isc/common/medm/OAF_OVERVIEW_BAK.adl
M       isc/common/medm/OAF_ADAPTATION_MICH.adl
controls@pianosa:/opt/rtcds/userapps/release 0$ 

This doesn't even include things that haven't even been added yet.  It doesn't take much time.  Just copy and paste what you elog about the changes.

  6721   Wed May 30 22:51:32 2012 DenUpdatePEMguralp isolation box

When I've put Guralps inside the isolation box, the signal from seismometers increased and was out of AA board range. I've reduced the gain of the readout box by a factor of 2. Now R2 for channels 1-6 is (2000, 1050, 1050, 2000, 1050, 1050) Ohm.

The signal increased in the frequency range 30-50 Hz. Guralp noise become better. That's good. However, it is still worse then in the manual.

As Yuta is dancing on the isolation box, Guralp signal is most time out of the AA board range. So I calculated the noise based on 5 min data. This may be enough, but I'll repeat the experiment later with 30 min data.

lin_box_noise.png

  6720   Wed May 30 22:04:13 2012 JenneBureaucracyGeneral40m Action Items

.....including:

fix Q of Vio2 filter in SUS. - JENNE

Switch power source for beatnote PD's amplifiers from temporary power supplies under PSL table to permanent taking the power from a rack.

  6719   Wed May 30 20:12:15 2012 ranaUpdateIOOMC beam spot oscillation

This is a common occurrence when diagnostic scripts are written without the ability to handle exceptions (e.g. ctrl-c, terminal gets closed, etc.).

The first thing to do is make sure that the "new" script you are writing doesn't already exist (hint: look in the old scripts directory).

If you are writing a script that touches things in the interferometer, it must always return the settings to the initial state on abnormal termination:

http://linuxdevcenter.com/pub/a/linux/lpt/44_12.html

  6718   Wed May 30 19:27:38 2012 yutaUpdateIOOMC beam spot oscillation

[Koji, Yuta]

We found that C1:SUS-MC{1,2,3}_TO_COIL_3_4_GAIN was somehow changed to -1, and feedback signal for SIDE was fedback to LLCOIL, which is apparently not correct.
We checked the snapshots on May 25 and confirmed that it was used to be 0, so we fixed it.
We suspect that it happened during the beam spot measurement, because the measurement changes the TO_COIL matrix gains.

Now, we don't see any MC beam spot oscillation.

Quote:

[Koji, Suresh, Jenne, Yuta]

Background:
  We noticed that the beam spots on MC mirrors are oscillating in ~ 1 Hz yesterday. It means MC mirrors are actually oscillating. This was observable even if the WFS servo is off.

 

  6717   Wed May 30 18:16:44 2012 JamieUpdateLSCskeleton of new c1cal calibration model created

[Jamie, Xavi Siemens, Chris Pankow]

We built the skeleton of a new calibration model for the LSC degrees of freedom.  I named it "c1cal".  It will run on the c1lsc FE machine, in CPU slot 4, and has been given DCUID 50.

Right now there's not much in the model, just inputs for DARM_ERR and DARM_CTRL, filters for each input, and the sum of the two channels that is h(t).

Tomorrow we'll extract all the needed signals from c1lsc, and see if we can generate something resembling a calibrated signal for one of the IFO DOFs.

 

  6716   Wed May 30 18:08:40 2012 JamieUpdateLSCc1lsc: add error point pick-offs, moved ctrl pick-offs after feedforward

I made some modifications to the c1lsc model in order to extract both the error and control signals.

I added pick-offs for the error signals right before IFO DOF filter modules.  These are then sent with GOTOs to outputs.

I also modified things on the control side.  The OAF stuff was picking off control signals before feedforward in/outs.  After discussing with Jenne we decided that it would make sense for the OAF to be looking at the control signals after feedforward.  It also makes sense to define the control signal after the feedforward.  These control signals are then sent with GOTOs to another set of outputs.

Finally, I moved the triggers to after the control signal pickoffs, and right before the output matrix.  The final chain looks like (see attachment):

input matrix --> power norm --> ERR pickoff --> DOF filters --> FF out --> FF in --> CTRL pickoff --> trigger --> output matrix

The error pickoff outputs in the top level of the model are left terminated for the moment.  Eventually I will be hooking these into the new c1cal calibration model.

The model was recompiled, installed, and restarted.  Everything came up fine.

Attachment 1: LSCchain.png
LSCchain.png
  6715   Wed May 30 15:51:22 2012 yutaUpdateIOOMC beam spot oscillation

[Koji, Suresh, Jenne, Yuta]

Background:
  We noticed that the beam spots on MC mirrors are oscillating in ~ 1 Hz yesterday. It means MC mirrors are actually oscillating. This was observable even if the WFS servo is off.

What we did:
  1. By measuring the spectra of OSEM sensor outputs, we found that MC3 is the one that is oscillating.

  2.  Oscillation at ~ 1 Hz only happened when the local damping using OSEMs are on (see Attachment 1; REF is when the damping is on).

  3.  We found that this oscillation came from insufficiency in phase margin in SUSPOS loop. So, we increased the gain, C1:SUS-MC3_SUSPOS_GAIN, from 95 to 200. It helped a little, but oscillation is still there.

  4.  We measured openloop transferfunctions of SUSPOS, SUSPIT, SUSYAW, SUSSIDE loop, and concluded that diagonalization some how went wrong. The amplitude of the oscillation (peak height in the OSEM spectra) changed by pushing the MC SUS connectors.

Plan:
  - Fix the connectors so that we don't have to push them any more.
  - Redo the diagonalization of the MC suspensions.

Attachment 1: specMC3_onoff_localdamping.pdf
specMC3_onoff_localdamping.pdf
  6714   Wed May 30 13:24:08 2012 JenneUpdateLockingYarm error spectra

Quote:

The ~16Hz bounce mode of some optic is showing up in the Yarm error signal. 

MC is kind of 'windy' looking, so maybe it's from that? (Yuta's guess).

We need to make sure that the SUS damping and oplev paths both have notches at the correct bounce mode, not the old, old MOS frequency.  If that doesn't work, may need to put a resgain in Yarm path.

 Made the Bounce notch in the BounceRoll filter (ITMY OLPIT, ITMY OLYAW)  wider, so it actually spans the peak we see in the error spectra.  When we next lock the arm later today, I'll retake this spectra to see if the ETMY oplev fix (Koji, Yuta) and this notch fix both helped.

  6713   Wed May 30 01:35:15 2012 yutaUpdateGreen Lockingaligned Y arm green beam

[Jenne, Yuta]

We aligned the Y arm for IR (C1:LSC-TRY_OUT is now ~ 0.9), and aligned the green beam from the ETMY table. The Y arm green is now resonating in TEM00 mode, but we need some monitors (green trans or green refl) to maximize the coupling.

We noticed that the MC beam spot are oscillating at ~ 1 Hz, mostly in YAW.  This wasn't observable before the PMC realignment (elog #6708). We should find out why and fix it.

  6712   Tue May 29 22:48:37 2012 DenUpdatePEMGuralp noise

I've checked whether the Guralp noise that we see comes not from seismometer but from ADC or readout box. I did 2 separate measurements . First, I've split 1 signal from Guralp into 2 before the input to AA board and subtracted one from another using Wiener filter. Second, I've connected inputs of channels 1 and 4 of the seismometer readout box and put the signal from seismometer to channel 1.

split_noises.png

The plot shows that ADC and readout box do not contribute too much to the Guralp noise.

  6711   Tue May 29 21:50:21 2012 JenneUpdateLockingYarm error spectra

The ~16Hz bounce mode of some optic is showing up in the Yarm error signal. 

MC is kind of 'windy' looking, so maybe it's from that? (Yuta's guess).

We need to make sure that the SUS damping and oplev paths both have notches at the correct bounce mode, not the old, old MOS frequency.  If that doesn't work, may need to put a resgain in Yarm path.

Attachment 1: LSC_POY_11_I_ERR_29May2012.pdf
LSC_POY_11_I_ERR_29May2012.pdf
  6710   Tue May 29 21:34:08 2012 JenneUpdateSUSETMY oplev spot size reduced

Quote:

Quote:

Quote:

 

 The typical sign of a dying gas laser is that it glows for a few minutes only. The power supplies are fine.

Two new  JDS - Uniphase 1103P lasers ( NT64-104 )  arriving on Monday, May 21

 Yesterday I swapped in new He/Ne laser with output power 3.5 mW  The return spot on qpd is large ~6mm in diameter and 20,500 counts

The spot size reduction require similar layout as ETMX oplev.

 The oplev path is relayed and the spot size on the qpd is reduced. I still have to clean up and replace "Miki Mouse" lens holder.

 Flipped the sign on the ETMY oplev servo gain, since it was wrong.  (It was "-" for both, now it is "+" for both)

  6709   Tue May 29 21:05:30 2012 yutaUpdateIOOPMC, MC alignment are shit

Quote:

  [Keiko, Jenne]

PMC aligned.  Suresh is fixing the measure MC spot positions script, then we'll remeasure MC spot positions.

 [Suresh, Jenne, Yuta]

We measured the MC spot positions twice tonight. Procedure for measuring them is in elog #6688.
The results were;

spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
    3.3359    3.9595    2.3171   -7.7424   -0.8406    6.4884

spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
    3.2681    4.0052    2.2808   -7.3965   -0.7624    7.1302

The spot moved by about 0.5 mm since May 25, but we concluded that this displacement is negligible and difficult to be fixed by aligning PSL beam.

We'll align Y arm and X arm next.

  6708   Tue May 29 19:50:01 2012 JenneUpdateIOOPMC, MC alignment are shit

Quote:

The followings are a kind of daily check. Do this without any notice:

- Align PMC.

Quote:

[Yuta, Jenne]

PMC and MC alignment are both shit, although with the WFS on, the MC is pretty good.  We're leaving it for now, so that (a) we don't mess up Koji's work, and (b) we can work on the Xarm.  Steve is doing Yarm oplev stuff, so we'll do Yarm later.

 

 [Keiko, Jenne]

PMC aligned.  Suresh is fixing the measure MC spot positions script, then we'll remeasure MC spot positions.

  6707   Tue May 29 17:40:45 2012 JenneUpdateIOOPMC, MC alignment are shit

Quote:

[Yuta, Jenne]

PMC and MC alignment are both shit, although with the WFS on, the MC is pretty good.  We're leaving it for now, so that (a) we don't mess up Koji's work, and (b) we can work on the Xarm.  Steve is doing Yarm oplev stuff, so we'll do Yarm later.

 [Yuta, Jenne, Suresh]

We pushed on the MC SUS connectors at the back of the rack, and that helped bring MC3 back to where it should be.  Then we looked at MC RFPD DC, and adjusted the optics with the WFS off, so that the refl is ~0.56.  Then when we turn the WFS on, the alignment doesn't really change, so we have offloaded the WFS. 

Now we're measuring the spot positions to check where the MC is.  Then we'll align the arms, and align the green to the arms.

  6706   Tue May 29 15:55:22 2012 steveUpdateSUSETMY oplev spot size reduced

Quote:

Quote:

 

 The typical sign of a dying gas laser is that it glows for a few minutes only. The power supplies are fine.

Two new  JDS - Uniphase 1103P lasers ( NT64-104 )  arriving on Monday, May 21

 Yesterday I swapped in new He/Ne laser with output power 3.5 mW  The return spot on qpd is large ~6mm in diameter and 20,500 counts

The spot size reduction require similar layout as ETMX oplev.

 The oplev path is relayed and the spot size on the qpd is reduced. I still have to clean up and replace "Miki Mouse" lens holder.

There was no IP-ANG coming out of the chamber at this time!

Attachment 1: ETMYoplev.jpg
ETMYoplev.jpg
Attachment 2: IMG_1243.JPG
IMG_1243.JPG
  6705   Tue May 29 15:49:02 2012 KojiUpdateLockingETMX trans camera

Quote:

...will be helpful for acquiring lock after the vent.  We should install a camera at ETMX.

 Do that.

  6704   Tue May 29 15:48:31 2012 KojiUpdateIOOPMC, MC alignment are shit

The followings are a kind of daily check. Do this without any notice:

- Align PMC.

- Check MC spot position with the script (where is it located?). Ignore MC2 result as it can be arbitrary set.

- If the MC1/MC3 spots have moved it means that the PSL beam has moved. If the beam has moved, we should have some discussion what we should do.

- If the spot positions are about the same as before, align the MC mirrors. This should be done by nulling the WFS feedback. (Someone should make a simple script for this WFS offloading.)

------------

Then, start locking both arms

Quote:

[Yuta, Jenne]

PMC and MC alignment are both shit, although with the WFS on, the MC is pretty good.  We're leaving it for now, so that (a) we don't mess up Koji's work, and (b) we can work on the Xarm.  Steve is doing Yarm oplev stuff, so we'll do Yarm later.

 

  6703   Tue May 29 15:29:16 2012 JamieUpdateComputerslatest pynds installed on all new control room machines

The DASWG lscsoft package repositories have a lot of useful analysis software.  It is all maintained for Debian "sqeeze", but it's mostly installable without modification on Ubuntu 10.04 "lucid" (which is based on Debian squeeze).  Basically the only thing that needs to access the lscsoft repositories is to add the following repository file:

controls@rossa:~ 0$ cat /etc/apt/sources.list.d/lscsoft.list 
deb http://www.lsc-group.phys.uwm.edu/daswg/download/software/debian/ squeeze contrib
deb-src http://www.lsc-group.phys.uwm.edu/daswg/download/software/debian/ squeeze contrib

deb http://www.lsc-group.phys.uwm.edu/daswg/download/software/debian/ squeeze-proposed contrib
deb-src http://www.lsc-group.phys.uwm.edu/daswg/download/software/debian/ squeeze-proposed contrib
controls@rossa:~ 0$ 

A simple "apt-get update" then makes all the lscsoft packages available.

lscsoft includes the nds2 client packages (nds2-client-lib) and pynds (python-pynds).  Unfortunately the python-pynds debian squeeze package currently depends on libboost-python1.42, which is not available in Ubuntu lucid.  Fortunately, pynds itself does not require the latest version and can use what's in lucid.  I therefore rebuilt the pynds package on one of the control room machines:

$ apt-get install dpkg-dev devscripts debhelper            # these are packages needed to build a debian/ubuntu package
$ apt-get source python-pynds                              # this downloads the source of the package, and prepares it for a package build
$ cd python-pynds-0.7
$ debuild -uc -us                                          # this actually builds the package
$ ls -al ../python-pynds_0.7-lscsoft1+squeeze1_amd64.deb
-rw-r--r-- 1 controls controls 69210 2012-05-29 11:57 python-pynds_0.7-lscsoft1+squeeze1_amd64.deb

I then copied the package into a common place:

/ligo/apps/debs/python-pynds_0.7-lscsoft1+squeeze1_amd64.deb

I then installed it on all the control room machines as such:

$ sudo apt-get install libboost-python1.40.0 nds2-client-lib python-numpy   # these are the dependencies of python-pynds
$ sudo dpkg -i /ligo/apps/debs/python-pynds_0.7-lscsoft1+squeeze1_amd64.deb

I did this on all the control room machines.

It looks like the next version of pynds won't require us to jump through these extra hoops and should "just work".

  6702   Tue May 29 14:59:39 2012 JenneUpdateIOOPMC, MC alignment are shit

[Yuta, Jenne]

PMC and MC alignment are both shit, although with the WFS on, the MC is pretty good.  We're leaving it for now, so that (a) we don't mess up Koji's work, and (b) we can work on the Xarm.  Steve is doing Yarm oplev stuff, so we'll do Yarm later.

  6701   Tue May 29 14:58:28 2012 JenneUpdateLockingETMX trans camera

...will be helpful for acquiring lock after the vent.  We should install a camera at ETMX.

ELOG V3.1.3-